Particles each with mass m and charge q accelerate through a potential difference of V = 100 volts and enter a region of uniform magnetic field B, where their paths are circular with radius R = 0.05 m as shown in Fig. 9.12. The magnitude of the field is B = 6.75 x 10βˆ’4 T. a. Indicate the direction of B on a drawing. b. Derive the following relationship between the radius R and q, m, V, and B: 𝑅 2 = π‘š π‘ž 2𝑉 𝐡2 c. Find the value of the charge-to-mass ratio (q/m) of these particles and identify them. Justify your answer! F


1
Expert's answer
2021-10-12T12:33:32-0400

Equate the forces acting on the particles: Lorentz force equals the centripetal force, then express r.


"qvB=m\frac{v^2}r,\\\space\\
r=\frac mq·\frac{v}{B}."

Express the velocity in terms of voltage and charge using energy conservation:


"v=\sqrt{\frac{2qV}{m}}."


Hence:


"r=\sqrt{\frac{m}{q}·\frac{2V}{B^2}},\\\space\\
r^2=\frac{m}{q}·\frac{2V}{B^2}."


The value of the charge-to-mass is


"\frac qm=\frac{2V}{(rB)^2}=175.583·10^9\text{ C/kg},"

which corresponds to charge-to-mass ratio for electron. You can check it by diving the charge of electron by its mass.